# A bi-fidelity method for the uncertain Vlasov-Poisson system near quasineutrality in an asymptotic-preserving PIC framework

### v1

@@ 5.3 The RF-VPME system: uncertain problem
5.3.1 Test 1: uncertain initial conditions
@@
In Fig. 7, we consider the model under different regimes with $\varepsilon=1,0.1,0.01$. This figure shows the errors of density $n$ and velocity $u$ between BF and HF solutions, as the number of HF simulation runs $r$ increases. A fast exponential decay of the errors is clearly observed in both plots, for different cases of $\varepsilon$. In addition, we can detect that the errors for these macroscopic quantities saturate as $r$ reaches about 10, at the accuracy level of $O(10^{-3})$. In Fig. 8, we test that even for large $\varepsilon$ ($\varepsilon=1$), the mean and standard deviation of BF approximations match really well with the HF solutions, by running the HF solver only 10 times.
In Fig. 9, we consider the cases $\varepsilon=1,0.1,0.01$ and compare the HF, LF and BF solutions of $n$ and $u$ at an arbitrarily selected $z$ point, by using $r=10$ in the BF approximation. It can be seen that the HF and BF solutions are consistent with each other, whereas the LF solutions display a relatively large difference at many spatial locations. Together with Fig. 8, these results suggest that even though our LF Euler model may not be accurate in the physical space for each fixed $z$, it still can capture the behavior and characteristics of the HF model (VPME system) in the random space, thus demonstrating the effectiveness and accuracy of our proposed BF approach.
Figure 7: (UQ Test 1, different $\varepsilon$ in Sec. 5.3.1) Errors of the bi-fidelity approximation for $n$ (left), $u$ (right) with respect to the number of high-fidelity simulation runs.
Figure 8: (UQ Test 1, $\varepsilon=1$ in Sec. 5.3.1) Mean and standard deviation of $n$, $u$ of high-fidelity and bi-fidelity solutions by using $r=10$.
@@
Figure 9: (UQ Test 1, different $\varepsilon$ in Sec. 5.3.1) Comparison of low-, high- and bi-fidelity solutions at a fixed $z$ by using $r=10$ in the bi-fidelity approximations.
5.3.2 Test 2: uncertain initial data in the double-delta form
@@
Here the random variable $\mathbf{z}=(z_{1},\cdots,z_{5})$ is 5-dimensional, with $\{z_{i}\}_{i=1}^{5}$ following the uniform distribution on $[-1,1]$. We set the final time $t=0.5$. In the HF model, $N_{h}=400$, $N_{tol}=1.6\times 10^{6}$ and $N_{c}=600$ is used, with the time step the same as Test 1. The numerical parameters for the LF model are the same as used in Test 1. We notice a fast decay of $L^{2}$ errors between HF and BF solutions in Fig. 10, and the errors achieve a satisfactory level of accuracy around $O(10^{-3})$ when the number of HF runs is as small as $r=5$. This phenomenon stays uniformly true for problem in different regimes, with $\varepsilon$ ranging from 0.01, 0.1 to 1. In Fig. 11, we test the model with $\varepsilon=0.01$ and show the mean and standard deviation of $n$ and $u$ by running the expensive HF solver only 10 times. One can observe that the BF approximations match really well with the HF solutions. Thus for our test 2 with more complex initial distribution, the same conclusion can be drawn as Test 1, that is, the BF model can effectively approximate the HF solution in different regimes, at a low computational cost.
Figure 10: (UQ Test 2 in Sec. 5.3.2) Errors of the bi-fidelity approximation for $n$ (left), $u$ (right) with respect to the number of high-fidelity simulation runs.
Figure 11: (UQ Test 2, $\varepsilon=0.01$ in Sec. 5.3.2) Mean and standard deviation of $n$, $u$ of high-fidelity and bi-fidelity solutions by using $r=10$.
5.3.3 Test 3: uncertain initial data with KL-expansion
@@
with $\tilde{n}(x)=\frac{1}{s}\left(1.5+0.2\cos{(2x)}+0.1\cos{(4x)}\right)$, where $s$ is the normalization coefficient. The detailed expansion of (5.12) is provided in Appendix 7.2. We consider the initial distribution given by
@@
Let the spatial domain $\mathcal{X}=[0,2\pi]$, the parameters in (7.8) be $\sigma=\frac{1}{15}$, $l=0.5$ with eigenvalues satisfying $\lambda_{j}>10^{-6}$. For the HF model, $N_{h}=400$, $\Delta t_{h}=0.4\frac{\Delta x}{\max\{u_{0}\}}$, $N_{tol}=1.6\times 10^{6}$ and $N_{c}=200$ repetition times is used to reduce the noise of PIC method. In the LF solver, we take $N_{l}=1000$ and $\Delta t_{l}=0.1\Delta x$. The model in different regimes ($\varepsilon=1,0.1,0.01$) is studied, with final time $t=0.5$.
In Fig. 12, we detect a simlar convergence trend as in the previous two tests, however the $l_{2}$ errors of the BF approximation decay slightly slower. For the model in different regimes ($\varepsilon=1,0.1,0.01$), the errors seem to be saturated when $r$ reaches about 30, at the accuracy level of nearly $O(10^{-4})$ for the density $n$ and $O(10^{-3})$ for the velocity. In this test, we employ the KL expansion to characterize the initial distribution that contains uncertainties, which is more complicated than the settings in previous two tests. As a result, for the BF construction one needs more information to capture the HF solution space, that is, through utilizing larger number of HF implementations one is able to approximate the HF model at the same accuracy level as in the previous tests.
Moreover, in Fig. 13 we present the expectation (left) and standard deviation (right) for quantities $n$, $u$ by using $r=30$ HF simulation runs in the BF approximation. It is clearly seen that the HF and BF solutions match quite well, thus demonstrating the efficiency and accuracy of our bi-fidelity method.
Figure 12: (UQ Test 3, $\varepsilon=1,0.1,0.01$ in Sec. 5.3.3) Errors of the bi-fidelity approximation for $n$ (left), $u$ (right) with respect to the number of high-fidelity simulation runs.
Figure 13: (UQ Test 3, $\varepsilon=0.1$ in Sec. 5.3.3) Mean and standard deviation of $n$, $u$ of high-fidelity solutions and bi-fidelity solutions by using $r=30$.
6 Conclusion
@@
Author contributions
Conceptualization: Liu Liu, Yanli Wang; Methodology: Guangwei Liu, Liu Liu, Yanli Wang; Formal analysis and investigation: Liu Liu, Yanli Wang; Writing - original draft preparation: Liu Liu, Yanli Wang; Software: Guangwei Liu, Liu Liu, Yanli Wang.
6.1 Acknowledgements
We thank Dr. Yiwen Lin (Shanghai Jiao Tong University) on helpful discussions for the Karhunen-Loéve expansions in our last numerical test.
Funding
L. Liu acknowledges the support by National Key R&D Program of China (2021YFA1001200), Ministry of Science and Technology in China, Early Career Scheme (24301021) and General Research Fund (14303022 & 14301423) funded by Research Grants Council of Hong Kong. This work of Yanli Wang is partially supported by the Foundation of the President of China Academy of Engineering Physics (YZJJZQ2022017), and the National Natural Science Foundation of China (Grant No. 12171026, U2230402, and 12031013).
Data availability
The datasets generated during and/or analyzed during the current study are available from the corresponding author on reasonable request.
Declarations
Competing interests The authors have no relevant financial or non-financial interests to disclose.
7 Appendix
7.1 Details of classical PIC method
@@
$n^{0}_{j}$ $=\frac{1}{h}\int_{x_{j}-\frac{h}{2}}^{x_{j}+\frac{h}{2}}\mathrm{d}x\>\int_{-\infty}^{+\infty}f^{0}(x,v)\mathrm{d}v$ (7.1)
@@
$W(x)=\left\{\begin{array}[]{ll}1&|x|<h/2\text{ or }x=h/2,\\[4.0pt] 0&\text{ otherwise }\end{array}\right.$ (7.2)
@@
$\frac{1}{h}\sum_{\begin{subarray}{c}\text{ particles k}\\ \text{ in cell }j\end{subarray}}w_{k}=n^{0}_{j}.$ (7.3)
For convenience, the initial number of particles per cell is set as the same, which depends on the total number of particles $N_{\text{total}}$ and the size of the mesh $N_{h}$. Besides, the weights of particles in same cell are set as the same initially, which means for all particles in cell $j$ at $t=0$, their weights can be derived by (7.3)
$w_{k}=\frac{hn^{0}_{j}}{M},\qquad M=\frac{N_{\text{total}}}{N_{h}}.$ (7.4)
@@
Theorem 1.
@@
$U=\sum_{j\in\mathbb{N}}Z_{j}\varphi_{j}.$ (7.5)
$Z_{j}=\int_{\mathcal{X}}U(x)\varphi_{j}(x)\mathrm{d}x.$ (7.6)
$\mathbb{E}_{\mu}\left[Z_{j}\right]=0,\qquad\mathbb{E}_{\mu}\left[Z_{j}Z_{k}\right]=\lambda_{j}\delta_{jk}.$ (7.7)
[…]
@@
$\varphi_{j}(x)=\left\{\begin{array}[]{ll}\sqrt{\frac{1}{\Lambda}},&j=0\\ \sqrt{\frac{2}{\Lambda}}\cos\left(\frac{2j\pi x}{\Lambda}\right),&j>1,\text{ even }\\ \sqrt{\frac{2}{\Lambda}}\sin\left(\frac{2j\pi x}{\Lambda}\right),&j>1,\text{ odd }\end{array}\right.$ (7.9)
@@
$\lambda_{j}=\int_{-\pi}^{\pi}c(x)\varphi_{j}(x)\mathrm{d}x.$ (7.10)
and $\{z_{j}\}_{j=0}^{\infty},\>z_{j}\sim\mathcal{N}(0,1)$. Note that the orthonormal basis (7.9) is not unique. Here, we choose the trigonometric function system.

### v2

@@ 5.3. The RF-VPME system: uncertain problem
5.3.1. Test 1: uncertain initial conditions
@@
In Figure 7, we consider the model under different regimes with $\varepsilon=1,0.1,0.01$. This figure shows the errors of density $n$ and velocity $u$ between BF and HF solutions, as the number of HF simulation runs $r$ increases. A fast exponential decay of the errors is clearly observed in both plots, for different cases of $\varepsilon$. In addition, we can detect that the errors for these macroscopic quantities saturate as $r$ reaches about 10, at the accuracy level of $O(10^{-3})$. In Figure 8, we test that even for large $\varepsilon$ ($\varepsilon=1$), the mean and standard deviation of BF approximations match really well with the HF solutions, by running the HF solver only 10 times.
In Figure 9, we consider the cases $\varepsilon=1,0.1,0.01$ and compare the HF, LF and BF solutions of $n$ and $u$ at an arbitrarily selected $z$ point, by using $r=10$ in the BF approximation. It can be seen that the HF and BF solutions are consistent with each other, whereas the LF solutions display a relatively large difference at many spatial locations. Together with Figure 8, these results suggest that even though our LF Euler model may not be accurate in the physical space for each fixed $z$, it still can capture the behavior and characteristics of the HF model (VPME system) in the random space, thus demonstrating the effectiveness and accuracy of our proposed BF approach.
Figure 7. (UQ Test 1, different $\varepsilon$ in Subsection 5.3.1) Errors of the bi-fidelity approximation for $n$ (left), $u$ (right) with respect to the number of high-fidelity simulation runs.
Figure 8. (UQ Test 1, $\varepsilon=1$ in Subsection5.3.1) Mean and standard deviation of $n$, $u$ of high-fidelity and bi-fidelity solutions by using $r=10$.
@@
Figure 9. (UQ Test 1, different $\varepsilon$ in Subsection 5.3.1) Comparison of low-, high- and bi-fidelity solutions at a fixed $z$ by using $r=10$ in the bi-fidelity approximations.
5.3.2. Test 2: uncertain initial data in the double-delta form
@@
Here the random variable $\mathbf{z}=(z_{1},\cdots,z_{5})$ is 5-dimensional, with $\{z_{i}\}_{i=1}^{5}$ following the uniform distribution on [1,1][-1,1]. We set the final time $t=0.5$. In the HF model, $N_{h}=400$, $N_{tol}=1.6\times 10^{6}$ and $N_{c}=600$ is used, with the time step the same as Test 1. The numerical parameters for the LF model are the same as used in Test 1. We notice a fast decay of $L^{2}$ errors between HF and BF solutions in Figure 10, and the errors achieve a satisfactory level of accuracy around $O(10^{-3})$ when the number of HF runs is as small as $r=5$. This phenomenon stays uniformly true for problem in different regimes, with $\varepsilon$ ranging from 0.01, 0.1 to 1. In Figure 11, we test the model with $\varepsilon=0.01$ and show the mean and standard deviation of $n$ and $u$ by running the expensive HF solver only 10 times. One can observe that the BF approximations match really well with the HF solutions. Thus for our test 2 with more complex initial distribution, the same conclusion can be drawn as Test 1, that is, the BF model can effectively approximate the HF solution in different regimes, at a low computational cost.
Figure 10. (UQ Test 2 in Subsection 5.3.2) Errors of the bi-fidelity approximation for $n$ (left), $u$ (right) with respect to the number of high-fidelity simulation runs.
Figure 11. (UQ Test 2, $\varepsilon=0.01$ in Subsection 5.3.2) Mean and standard deviation of $n$, $u$ of high-fidelity and bi-fidelity solutions by using $r=10$.
5.3.3. Test 3: uncertain initial data with KL-expansion
@@
with $\tilde{n}(x)=\frac{1}{s}\left(1.5+0.2\cos{(2x)}+0.1\cos{(4x)}\right)$, where $s$ is the normalization coefficient. The detailed expansion of (5.12) is provided in Appendix A.3. We consider the initial distribution given by
@@
Let the spatial domain $\mathcal{X}=[0,2\pi]$, the parameters in (A.12) be $\sigma=\frac{1}{15}$, $l=0.5$ with eigenvalues satisfying $\lambda_{j}>10^{-6}$. For the HF model, $N_{h}=400$, $\Delta t_{h}=0.4\frac{\Delta x}{\max\{u_{0}\}}$, $N_{tol}=1.6\times 10^{6}$ and $N_{c}=200$ repetition times is used to reduce the noise of PIC method. In the LF solver, we take $N_{l}=1000$ and $\Delta t_{l}=0.1\Delta x$. The model in different regimes ($\varepsilon=1,0.1,0.01$) is studied, with final time $t=0.5$.
In Figure 12, we detect a simlar convergence trend as in the previous two tests, however the $l_{2}$ errors of the BF approximation decay slightly slower. For the model in different regimes ($\varepsilon=1,0.1,0.01$), the errors seem to be saturated when $r$ reaches about 30, at the accuracy level of nearly $O(10^{-4})$ for the density $n$ and $O(10^{-3})$ for the velocity. In this test, we employ the KL expansion to characterize the initial distribution that contains uncertainties, which is more complicated than the settings in previous two tests. As a result, for the BF construction one needs more information to capture the HF solution space, that is, through utilizing larger number of HF implementations one is able to approximate the HF model at the same accuracy level as in the previous tests.
Moreover, in Figure 13 we present the expectation (left) and standard deviation (right) for quantities $n$, $u$ by using $r=30$ HF simulation runs in the BF approximation. It is clearly seen that the HF and BF solutions match quite well, thus demonstrating the efficiency and accuracy of our bi-fidelity method.
Figure 12. (UQ Test 3, $\varepsilon=1,0.1,0.01$ in Figure 5.3.3) Errors of the bi-fidelity approximation for $n$ (left), $u$ (right) with respect to the number of high-fidelity simulation runs.
Figure 13. (UQ Test 3, $\varepsilon=0.1$ in Subsection 5.3.3) Mean and standard deviation of $n$, $u$ of high-fidelity solutions and bi-fidelity solutions by using $r=30$.
6. Conclusion
@@
Acknowledgements
{acknowledgement}
L. Liu acknowledges the support by National Key R&D Program of China (2021YFA1001200), Ministry of Science and Technology in China, Early Career Scheme (24301021) and General Research Fund (14303022 & 14301423) funded by Research Grants Council of Hong Kong. This work of Yanli Wang is partially supported by the Foundation of the President of China Academy of Engineering Physics (YZJJZQ2022017), and the National Natural Science Foundation of China (Grant No. 12171026, U2230402, and 12031013). We also thank Dr. Yiwen Lin (Shanghai Jiao Tong University) on helpful discussions for the Karhunen-Loéve expansions in our last numerical test.
[…]
A.1. Derivation of isothermal compressible Euler system
We gives the details of deriving the isothermal compressible Euler system given by equation (2.4) here. Taking the first and second moments in velocity for equation (2.1a), one gets the mass and momentum conservation equations given by
$\left\{\begin{array}[]{l}\partial_{t}n+\nabla_{\mathbf{x}}\cdot(n\mathbf{u})=0,\\[6.0pt] \partial_{t}(n\mathbf{u})+\nabla_{\mathbf{x}}\cdot(n\mathbf{u}\otimes\mathbf{u})+\nabla\!\cdot P=-\,n\nabla\phi.\end{array}\right$. (A.1)
Here $P$ is the non-dimensional ion pressure tensor. If we assume isotropy $P=pI$ (with $I$ being the identity matrix), then $\nabla\!\cdot P=\nabla p$. The usual procedure to close this system is to assume a thermodynamic equation that describes the relation between $p$ and $n$, which is shown as
$p=\tau n,\qquad\tau=\frac{T_{i}}{T_{e}}.$ (A.2)
This closure assumption for the ion moment system can be found in [39, Section 2]. Based on this closed system, one considers the Euler-Poisson-Boltzmann (EPB) model [47] by neglecting the ion pressure tensor, where the cold-ion approximation (i.e., $T_{i}\to 0$, so $\tau\ll 1$) is adopted:
$\left\{\begin{array}[]{l}\partial_{t}n+\nabla_{\mathbf{x}}\cdot(n\mathbf{u})=0,\\[6.0pt] \partial_{t}(n\mathbf{u})+\nabla_{\mathbf{x}}\cdot(n\mathbf{u}\otimes\mathbf{u})=-\,n\nabla\phi.\end{array}\right$. (A.3)
If the quasi-neutral assumption is made, the Poisson equation (2.1b) is replaced by the constraint of local charge being zero:
$n=e^{\phi}.$
Then one can write
$n\nabla\phi=e^{\phi}\nabla\phi=\nabla e^{\phi}=\nabla n,$
and the quasi-neutral Euler-Poisson-Boltzmann model is exactly the Isothermal Compressible Euler model given in equation (2.4):
$\left\{\begin{array}[]{l}\partial_{t}n+\nabla_{\mathbf{x}}\cdot(n\mathbf{u})=0,\\[6.0pt] \partial_{t}(n\mathbf{u})+\nabla_{\mathbf{x}}\cdot(n\mathbf{u}\otimes\mathbf{u})=-\nabla n.\end{array}\right$. (A.4)
A.2. Details of the classical PIC method
@@
$n^{0}_{j}$ $=\frac{1}{h}\int_{x_{j}-\frac{h}{2}}^{x_{j}+\frac{h}{2}}\mathrm{d}x\>\int_{-\infty}^{+\infty}f^{0}(x,v)\mathrm{d}v$ (A.5)
@@
$W(x)=\left\{\begin{array}[]{ll}1&|x|<h/2\text{ or }x=h/2,\\[4.0pt] 0&\text{ otherwise }\end{array}\right$. (A.6)
@@
$\frac{1}{h}\sum_{\begin{subarray}{c}\text{ particles k}\\ \text{ in cell }j\end{subarray}}w_{k}=n^{0}_{j}.$ (A.7)
For convenience, the initial number of particles per cell is set as the same, which depends on the total number of particles $N_{\text{total}}$ and the size of the mesh $N_{h}$. Besides, the weights of particles in same cell are set as the same initially, which means for all particles in cell $j$ at $t=0$, their weights can be derived by (A.7)
$w_{k}=\frac{hn^{0}_{j}}{M},\qquad M=\frac{N_{\text{total}}}{N_{h}}.$ (A.8)
@@
$U=\sum_{j\in\mathbb{N}}Z_{j}\varphi_{j}.$ (A.9)
$Z_{j}=\int_{\mathcal{X}}U(x)\varphi_{j}(x)\mathrm{d}x.$ (A.10)
$\mathbb{E}_{\mu}\left[Z_{j}\right]=0,\qquad\mathbb{E}_{\mu}\left[Z_{j}Z_{k}\right]=\lambda_{j}\delta_{jk}.$ (A.11)
@@
$c(x-y)=\sigma^{2}e^{-\frac{\left|x-y\right|^{2}}{l^{2}}},\>0<l\ll\Lambda.$ (A.12)
@@
$\varphi_{j}(x)=\left\{\begin{array}[]{ll}\sqrt{\frac{1}{\Lambda}},&j=0\\ \sqrt{\frac{2}{\Lambda}}\cos\left(\frac{2j\pi x}{\Lambda}\right),&j>1,\text{ even }\\ \sqrt{\frac{2}{\Lambda}}\sin\left(\frac{2j\pi x}{\Lambda}\right),&j>1,\text{ odd }\end{array}\right$. (A.13)
@@
$\lambda_{j}=\int_{-\pi}^{\pi}c(x)\varphi_{j}(x)\mathrm{d}x.$ (A.14)
and $\{z_{j}\}_{j=0}^{\infty},\>z_{j}\sim\mathcal{N}(0,1)$. Note that the orthonormal basis (A.13) is not unique. Here, we choose the trigonometric function system.